OpenStudy (darkbluechocobo):

In what distance will a car traveling with 100 kN stop if a force of -200 N is applied

OpenStudy (darkbluechocobo):

@Data_LG2 What would the equation look like for this?

OpenStudy (anonymous):

just to make sure... kN is kiloNewton right?

OpenStudy (darkbluechocobo):

Yes

OpenStudy (anonymous):

hmm... give me a sec to analyse this one.. "car traveling with 100 kN stop if a force of -200 N is applied" this part doesn't make sense to me

OpenStudy (darkbluechocobo):

Yes this problem seems strange and I dont know which equation to use s:

OpenStudy (anonymous):

Probably we can use kN to calculate for the mass using F=mg,, F=100 kN and g=9.8 m/s^2 Then use this to calculate for the acceleration using F= -200 N The car went stop so Vf=0m/s ....we are looking for the distance traveled.. hmm...

OpenStudy (darkbluechocobo):

so 100*9.8 F=980 ?

OpenStudy (anonymous):

m= F/g

OpenStudy (anonymous):

brb

OpenStudy (darkbluechocobo):

Alright :p sorry I'm still kind of new to stuff like this S:

OpenStudy (anonymous):

nvm, it still doesn't make sense :c maybe @aaronq can help

OpenStudy (darkbluechocobo):

man s:evil question

OpenStudy (darkbluechocobo):

hmm could you help find equation for How much force is applied to stop a 1,000 kg car moving 20m/s? the distance is 10m

OpenStudy (darkbluechocobo):

a lot more info is used

OpenStudy (anonymous):

\(\sf V_i = 20m/s, V_f=0m/s,\ m=1000 kg, d=10 m\) Use the formulas that we used before \(\sf V_f^2=V_i^2 +2ad\) use this to solve for acceleration, then use \(\sf F=ma\) to solve for the Force

OpenStudy (darkbluechocobo):

So we have to get A alone s: Like my weakness with equations is like turning them around to solve for what is needed so a=20^2+2(10)

OpenStudy (darkbluechocobo):

-V^2f at the end s:

OpenStudy (anonymous):

let's see... \(\sf \sf V_f^2=V_i^2 +2ad\\0=V_i^2+2ad\\-2ad=V_i^2\\a=\large \frac{V_i^2}{-2d}\)

OpenStudy (darkbluechocobo):

so basically what we did with the last one?

OpenStudy (darkbluechocobo):

wow i suck with equations s: so 400/-20

OpenStudy (darkbluechocobo):

so -20 m/s is the acceleration

OpenStudy (anonymous):

similar with that, but you were solving for \(\sf V_i\) before but now you're solving for a. yes, that's right

OpenStudy (darkbluechocobo):

1000*(-20)=-20000 N?

OpenStudy (darkbluechocobo):

so that would be le force

OpenStudy (anonymous):

yes... ^_^ lol I'm still thinking about the question that you posted ... i'll try to find out the solution, but if you find out how to do it, can you tell me? I want to know how to do thatxD

OpenStudy (darkbluechocobo):

Alright I shall tell you when I find out :p

OpenStudy (anonymous):

thanks!

OpenStudy (darkbluechocobo):

thanks for helping me ^-^

OpenStudy (darkbluechocobo):

Do you work a lot with physics?

OpenStudy (anonymous):

no not really, but I like the subject :P I hate it at first but then I keep solving and solving physics probs then I started to enjoy learning physics and math

OpenStudy (darkbluechocobo):

oh I see